OpenStudy (anonymous):

the sum of one half of a number and 6 equals one third of the number what is the number

ganeshie8 (ganeshie8):

let that number be x \(\huge \frac{x}{2} + 6 = \frac{x}{3}\) 3x + 36 = 2x x = 36

OpenStudy (jiteshmeghwal9):

Let the no. be 'x'. Then, its half will be:-\[{1 \over x} \times 6={x \over 6}\] & its one third will be:-\[\frac{1}{3} \times x=\frac{x}{3}\] A/q, \[\frac{x}{6}+6=\frac{x}{3}\]\[{x+36 \over 6}={x \over 3}\] Now, do cross multiplication as\[3(x+36)=6(x)\]\[3x+108=6x\]\[108=6x-3x\]\[108=3x\]\[{108 \over 3}=x\]\[36=x\]